# Rationale Matters: Learning Transferable Rubrics via Proxy-Guided Critique for VLM Reward Models

### 17 Mar 2026 v1

@@ Abstract
We propose Proxy-GRM, which introduces proxy-guided rubric verification into Reinforcement Learning (RL) to explicitly enhance rubric quality. Concretely, we train lightweight proxy agents (Proxy-SFT and Proxy-RL) that take a candidate rubric together with the original query and preference pair, and then predict the preference ordering using only the rubric as evidence. The proxy’s prediction accuracy serves as a rubric-quality reward, incentivizing the model to produce rubrics that are internally consistent and transferable. With ${\sim}50\text{k}$ data samples, Proxy-GRM reaches state-of-the-art results on the VL-Reward Bench, Multimodal Reward Bench, and MM-RLHF-Reward Bench, outperforming the methods trained on four times the data. Ablations show Proxy-SFT is a stronger verifier than Proxy-RL, and implicit reward aggregation performs best. Crucially, the learned rubrics transfer to unseen evaluators, improving reward accuracy at test time without additional training. Our code is available at https://github.com/Qwen-Applications/Proxy-GRM.
@@ 1 Introduction
  • We achieve state-of-the-art results on VL-RewardBench, Multimodal Reward Bench, and MM-RLHF-Reward Bench with only ${\sim}50\text{k}$ training samples (${\approx}4\times$ less than comparable methods), with rubrics that generalize to unseen evaluators without retraining.
@@ 2.1 Reward Models for Vision-Language Models
Reward models serve as a critical component in aligning large language models and vision-language models with human preferences (christiano2017deep; stiennon2020learning; bai2022training). Early approaches primarily employed discriminative reward models that output scalar scores (bradley1952rank; zhu2024starling). Bradley–Terry models (ouyang2022training) and their variants have been widely used to parameterize pairwise preferences. In the multimodal domain, models such as IXC-2.5-Reward (zang2025internlm), MM-RLHF-Reward (zhang2025mm), and LLaVA-Critic (xiong2025llava) have extended these ideas to handle image-text inputs.
@@ 4.1 Implementation Details
  • MM-RLHF-Reward Bench (zhang2025mm): A challenging benchmark including multiple-choice, long-form, short-form, safety, and video evaluation tasks.
Baselines. We compare against: (1) closed-source models including GPT-4o, Claude-3.5/3.7-Sonnet, and Gemini-1.5-Pro/Flash; (2) open-source reward models including IXC-2.5-Reward (zang2025internlm), R1-Reward (zhang2025r1), MM-RLHF-Reward (zhang2025mm), and Unified-Reward-Think (wang2025unifiedthink); and (3) our own ablation baselines Proxy-GRM-SFT (cold-start only) and Proxy-GRM-RL (RL without proxy reward).
Table 1: Combined Results on VL-RewardBench, Multimodal Reward Bench, and MM-RLHF-Reward Bench. Best results are marked with a green background and second best with a yellow background.
Model Proxy Agent Data VL-RewardBench Multimodal MM-RLHF-Reward Bench
@@ 4.2 Main Results
Results on MM-RLHF-Reward Bench. On this benchmark, Proxy-GRM-RL with Proxy-SFT achieves the highest Acc of 82.94%, surpassing R1-Reward (80.59%) and all closed-source models except Claude-3.7-Sonnet on the Acc+ metric. The progressive improvement from Proxy-GRM-SFT (79.41%) to Proxy-GRM-RL (82.94%) clearly demonstrates the complementary benefits of reinforcement learning and proxy-guided rubric verification.
@@
Proxy Agents Avg Score VL-Reward Bench Multimodal Reward Bench MM-RLHF- Reward Bench
@@ 4.3 Analysis Studies
Effect of model scale. Among general-purpose instruction models, larger models yield better proxy performance: Qwen2.5-VL-32B-Instruct (74.02% on VL-RewardBench) outperforms Qwen2.5-VL-7B-Instruct (73.70%), which in turn outperforms Qwen2.5-VL-3B-Instruct (72.25%). However, the 3B model actually degrades performance below the no-proxy baseline on MM-RLHF-Reward Bench (76.48% vs. 80.59%), indicating that insufficiently capable proxy agents introduce harmful noise into the training signal. As illustrated in Figure˜3, the 3B model fails to distinguish between high-quality and low-quality rubrics during early RL training, providing essentially random feedback that misleads the policy model about what constitutes a good rubric.
@@
Table 4: Rubric Transferability Results on VL-Reward-Bench, Multimodal Reward Bench, MM-RLHF-Reward Bench.
Evaluator Rubric Source VL-Reward Bench Multimodal Reward Bench MM-RLHF- Reward Bench
@@ D.2 Results on Multimodal Reward Bench
D.3 Results on MM-RLHF-Reward Bench
MM-RLHF-Reward Bench is a challenging benchmark covering Multiple-Choice (MCQ), Long-form, Short-form, Safety, and Video evaluation tasks. We report the standard accuracy (Acc) and the stricter Acc+ metric, which requires correctness on the harder subset of questions. The detailed results are shown in Table˜8.
@@ D.3 Results on MM-RLHF-Reward Bench
Table 8: Results on MM-RLHF-Reward Bench. Best results are marked with a green background and second best with a yellow background.
@@ E.2 Ablation on Proxy-SFT Training Data Size
Data Size $\textbf{VL-RB}\penalty 10000\ \textbf{Overall}_{\text{Acc}}$ $\textbf{VL-RB}\penalty 10000\ \textbf{Macro}_{\text{Acc}}$ MM-RB Overall MM-RLHF-RB Acc MM-RLHF-RB Acc+

### 18 Mar 2026 v2

@@ Abstract
We propose Proxy-GRM, which introduces proxy-guided rubric verification into Reinforcement Learning (RL) to explicitly enhance rubric quality. Concretely, we train lightweight proxy agents (Proxy-SFT and Proxy-RL) that take a candidate rubric together with the original query and preference pair, and then predict the preference ordering using only the rubric as evidence. The proxy’s prediction accuracy serves as a rubric-quality reward, incentivizing the model to produce rubrics that are internally consistent and transferable. With ${\sim}50\text{k}$ data samples, Proxy-GRM reaches state-of-the-art results on the VL-Reward Bench, Multimodal Reward Bench, and $M$-RLHF-Reward Bench, outperforming the methods trained on four times the data. Ablations show Proxy-SFT is a stronger verifier than Proxy-RL, and implicit reward aggregation performs best. Crucially, the learned rubrics transfer to unseen evaluators, improving reward accuracy at test time without additional training. Our code is available at https://github.com/Qwen-Applications/Proxy-GRM.
@@ 1 Introduction
  • We achieve state-of-the-art results on VL-RewardBench, Multimodal Reward Bench, and $M$-RLHF-Reward Bench with only ${\sim}50\text{k}$ training samples (${\approx}4\times$ less than comparable methods), with rubrics that generalize to unseen evaluators without retraining.
@@ 2.1 Reward Models for Vision-Language Models
Reward models serve as a critical component in aligning large language models and vision-language models with human preferences (christiano2017deep; stiennon2020learning; bai2022training). Early approaches primarily employed discriminative reward models that output scalar scores (bradley1952rank; zhu2024starling). Bradley–Terry models (ouyang2022training) and their variants have been widely used to parameterize pairwise preferences. In the multimodal domain, models such as IXC-2.5-Reward (zang2025internlm), $M$-RLHF-Reward (zhang2025mm), and LLaVA-Critic (xiong2025llava) have extended these ideas to handle image-text inputs.
@@ 4.1 Implementation Details
  • $M$-RLHF-Reward Bench (zhang2025mm): A challenging benchmark including multiple-choice, long-form, short-form, safety, and video evaluation tasks.
Baselines. We compare against: (1) closed-source models including GPT-4o, Claude-3.5/3.7-Sonnet, and Gemini-1.5-Pro/Flash; (2) open-source reward models including IXC-2.5-Reward (zang2025internlm), R1-Reward (zhang2025r1), $M$-RLHF-Reward (zhang2025mm), and Unified-Reward-Think (wang2025unifiedthink); and (3) our own ablation baselines Proxy-GRM-SFT (cold-start only) and Proxy-GRM-RL (RL without proxy reward).
Table 1: Combined Results on VL-RewardBench, Multimodal Reward Bench, and $M$-RLHF-Reward Bench. Best results are marked with a green background and second best with a yellow background.
Model Proxy Agent Data VL-RewardBench Multimodal $M$-RLHF-Reward Bench
@@ 4.2 Main Results
Results on $M$-RLHF-Reward Bench. On this benchmark, Proxy-GRM-RL with Proxy-SFT achieves the highest Acc of 82.94%, surpassing R1-Reward (80.59%) and all closed-source models except Claude-3.7-Sonnet on the Acc+ metric. The progressive improvement from Proxy-GRM-SFT (79.41%) to Proxy-GRM-RL (82.94%) clearly demonstrates the complementary benefits of reinforcement learning and proxy-guided rubric verification.
@@
Proxy Agents Avg Score VL-Reward Bench Multimodal Reward Bench $M$-RLHF- Reward Bench
@@ 4.3 Analysis Studies
Effect of model scale. Among general-purpose instruction models, larger models yield better proxy performance: Qwen2.5-VL-32B-Instruct (74.02% on VL-RewardBench) outperforms Qwen2.5-VL-7B-Instruct (73.70%), which in turn outperforms Qwen2.5-VL-3B-Instruct (72.25%). However, the 3B model actually degrades performance below the no-proxy baseline on $M$-RLHF-Reward Bench (76.48% vs. 80.59%), indicating that insufficiently capable proxy agents introduce harmful noise into the training signal. As illustrated in Figure˜3, the 3B model fails to distinguish between high-quality and low-quality rubrics during early RL training, providing essentially random feedback that misleads the policy model about what constitutes a good rubric.
@@
Table 4: Rubric Transferability Results on VL-Reward-Bench, Multimodal Reward Bench, $M$-RLHF-Reward Bench.
Evaluator Rubric Source VL-Reward Bench Multimodal Reward Bench $M$-RLHF- Reward Bench
@@ D.2 Results on Multimodal Reward Bench
D.3 Results on $M$-RLHF-Reward Bench
$M$-RLHF-Reward Bench is a challenging benchmark covering Multiple-Choice (MCQ), Long-form, Short-form, Safety, and Video evaluation tasks. We report the standard accuracy (Acc) and the stricter Acc+ metric, which requires correctness on the harder subset of questions. The detailed results are shown in Table˜8.
@@ D.3 Results on $M$-RLHF-Reward Bench
Table 8: Results on $M$-RLHF-Reward Bench. Best results are marked with a green background and second best with a yellow background.
@@ E.2 Ablation on Proxy-SFT Training Data Size
Data Size $\textbf{VL-RB}\penalty 10000\ \textbf{Overall}_{\text{Acc}}$ $\textbf{VL-RB}\penalty 10000\ \textbf{Macro}_{\text{Acc}}$ $M$-RB Overall $M$-RLHF-RB Acc $M$-RLHF-RB Acc+